OpenStudy (anonymous):

cot A - tan A = 2cot^2A

OpenStudy (anonymous):

with steps please !

OpenStudy (anonymous):

cotA=1/tanA 1/tanA - tan A =21/tan2a tan2a=1-tansuraea/1+tansquarea

OpenStudy (anonymous):

1/tanA - tan A =1-tan^2a/1+tan^2a 1-tan^2a/tana=1-tan^2a/1+tan^2a next u try

OpenStudy (anonymous):

okay tnx :)